# Achieving Linear Speedup with ProxSkip in Distributed Stochastic Optimization

### v3

@@ X Proof of Theorem 5
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}},$
@@
$\alpha\leq\min\left\{\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},\ \chi\geq\frac{288(1-p)}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}\leq\frac{1+\gamma}{2}<1.$
Since $\alpha\leq\frac{1-\lambda_{2}}{32\sqrt{3}\chi L}$ and $\frac{\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\leq\frac{8\chi^{2}}{(1-\lambda_{2})^{2}}$, we have $\frac{\alpha^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\leq\frac{1}{384L^{2}}$. Thus, it holds that
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{n\alpha^{2}L^{2}}{96}\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{n\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}.$
@@
$\left[\begin{array}[]{c}\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big{\|}^{2}\right]\\ \frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\right]\\ \end{array}\right]\leq\underbrace{\left[\begin{array}[]{cc}1-\mu\alpha&6\alpha L\\ \frac{\alpha^{2}L^{2}}{96}&\frac{1+\gamma}{2}\\ \end{array}\right]}_{:=A}\left[\begin{array}[]{c}\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big{\|}^{2}\right]\\ \frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}\right]\\ \end{array}\right]+\underbrace{\left[\begin{array}[]{c}\frac{\alpha^{2}\sigma^{2}}{n}\\ \frac{\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}\\ \end{array}\right]}_{:=b}.$ (77)
@@
$=\frac{1}{\mu\alpha(1-\gamma)(\frac{1}{2}-\frac{\alpha^{3}L^{3}}{16\mu(1-\gamma)})}\left[\begin{array}[]{cc}\frac{1-\gamma}{2}&6\alpha L\\ \frac{\alpha^{2}L^{2}}{96}&\mu\alpha\\ \end{array}\right]\left[\begin{array}[]{c}\frac{\alpha^{2}\sigma^{2}}{n}\\ \frac{\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}\\ \end{array}\right]$
$\leq\frac{4}{\alpha\mu(1-\gamma)}\left[\begin{array}[]{c}\frac{(1-\gamma)\alpha^{2}\sigma^{2}}{2n}+\frac{6L\alpha^{3}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}\\ \frac{\alpha^{4}L^{2}\sigma^{2}}{96n}+\frac{\mu\alpha^{3}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{192\chi^{2}}\\ \end{array}\right],$
@@
$\|(I-A)^{-1}b\|_{1}\leq\frac{2\alpha\sigma^{2}}{\mu n}+\frac{(6L\alpha^{2}\sigma^{2}+\mu\alpha^{2}\sigma^{2})(192\chi^{2}+(4\chi^{2}+2(1-p)))}{48\mu(1-\gamma)\chi^{2}}.$
@@
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big{\|}^{2}\right]\leq\Big{(}1-\frac{\alpha\mu}{4}\Big{)}^{t}a_{0}+\frac{2\alpha\sigma^{2}}{\mu n}+\frac{7L\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{48\mu(1-\gamma)\chi^{2}}.$
Since ${\bf{X}}^{0}=[{\bf{x}}^{0},\cdots,{\bf{x}}^{0}]^{\sf T}$, by [42, (75)], we have $\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq 2\alpha^{2}\|({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-1}\|\|\nabla F({\bf{X}}^{0})-{\bf{1}}_{n}\otimes(\nabla f({\bf{x}}^{0}))^{\sf T}\|^{2}=\frac{2n\alpha^{2}\varsigma_{0}^{2}}{1-\gamma}$. Note that $\frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}}$. It holds that
$a_{0}=\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}+\frac{1}{n}\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}+\frac{8\chi\alpha^{2}\varsigma_{0}^{2}}{1-\lambda_{2}}$
@@ XI Proof of Corollary 4
$c_{0}=\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2},\ b_{0}=\frac{8\chi\varsigma_{0}^{2}}{1-\lambda_{2}},\ c_{1}=\frac{2\sigma^{2}}{\mu n},\ c_{2}=\frac{7L\sigma^{2}(192\chi^{2}+(4\chi^{2}+2(1-p)))}{12\mu(1-\lambda_{2})\chi}.$
@@
$\alpha\leq\frac{1}{\underline{\alpha}}\triangleq\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\frac{72\mu}{L^{2}},\frac{1-\gamma}{12L+\nicefrac{{\mu}}{{2}}},\sqrt[3]{4\mu(1-\gamma)}\frac{1}{L}\right\}=\mathcal{O}\left(\frac{\mu(1-\lambda_{2})}{\chi L^{2}}\right).$
@@
Collecting these cases together into (XI), we have
@@ XII Proof of Lemma 5
$\widetilde{{\bf{X}}}^{t+1}$ $={\bf{W}}_{a}\widetilde{{\bf{Z}}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t},$ (80b)
$\widetilde{{\bf{U}}}^{t+1}$ $=\widetilde{{\bf{U}}}^{t}+\frac{p}{2\chi}{\bf{W}}_{b}\widetilde{{\bf{Z}}}^{t}+p{\bf{E}}^{t}.$ (80c)
@@
$\left[\begin{array}[]{c}\widetilde{{\bf{X}}}^{t+1}\\ \widetilde{{\bf{U}}}^{t+1}\\ \end{array}\right]=\left[\begin{array}[]{cc}{\bf{W}}_{a}&-{\bf{W}}_{a}{\bf{W}}_{b}\\ \frac{p}{2\chi}{\bf{W}}_{b}&{\bf{I}}-\frac{p}{2\chi}{\bf{W}}_{b}^{2}\\ \end{array}\right]\left[\begin{array}[]{c}\widetilde{{\bf{X}}}^{t}\\ \widetilde{{\bf{U}}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}{\bf{W}}_{a}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p}{2\chi}{\bf{W}}_{b}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]+\left[\begin{array}[]{c}-{\bf{W}}_{b}{\bf{E}}^{t}\\ p{\bf{E}}^{t}\\ \end{array}\right].$
@@
$\left[\begin{array}[]{c}{\bf{P}}^{-1}\widetilde{{\bf{X}}}^{t+1}\\ {\bf{P}}^{-1}\widetilde{{\bf{U}}}^{t+1}\\ \end{array}\right]=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{\Lambda}}}_{b}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}&{\bf{I}}-\frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]\left[\begin{array}[]{c}{\bf{P}}^{-1}\widetilde{{\bf{X}}}^{t}\\ {\bf{P}}^{-1}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}{\bf{P}}^{-1}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}{\bf{P}}^{-1}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]+\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}{\bf{P}}^{-1}{\bf{E}}^{t}\\ p{\bf{P}}^{-1}{\bf{E}}^{t}\\ \end{array}\right].$
@@
$\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t+1}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t+1}\\ \end{array}\right]$ $=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{\Lambda}}}_{b}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}&{\bf{I}}-\frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]+\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ p\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right].$
@@
${\bf{H}}^{\mathrm{s}}=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{\Lambda}}}_{b}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}&{\bf{I}}-\frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]=\left[\begin{array}[]{cc}{\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})&-({\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}}))\sqrt{{\bf{I}}-\hat{{\bf{\Lambda}}}}\\ \frac{p}{2\chi}\sqrt{{\bf{I}}-\hat{{\bf{\Lambda}}}}&{\bf{I}}-\frac{p}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})\\ \end{array}\right]$
@@
$H^{\mathrm{s}}_{i}=\left[\begin{array}[]{cc}1-\frac{1}{2\chi}(1-\lambda_{i})&-(1-\frac{1}{2\chi}(1-\lambda_{i}))\sqrt{1-\lambda_{i}}\\ \frac{p}{2\chi}\sqrt{1-\lambda_{i}}&1-\frac{p}{2\chi}(1-\lambda_{i})\\ \end{array}\right].$
Setting $\nu_{i}=1-\frac{1}{2\chi}(1-\lambda_{i})$, we have $\nu_{i}\in(0,1)$ and $H_{i}$ can be rewritten as
$H^{\mathrm{s}}_{i}=\left[\begin{array}[]{cc}\nu_{i}&-\nu_{i}\sqrt{2\chi(1-\nu_{i})}\\ \frac{p}{2\chi}\sqrt{2\chi(1-\nu_{i})}&1-p(1-\nu_{i})\\ \end{array}\right].$
@@
Since $\nu_{i}=1-\frac{1}{2\chi}(1-\lambda_{i})\geq 1-\frac{1}{2\chi}(1-\lambda_{n}),i=2,\ldots,n$ and $\lambda_{n}\in(-1,1)$, it holds that
$\chi\geq\frac{1}{p}\geq\frac{(1+p)^{2}}{4p}>\frac{(1-\lambda_{n})(1+p)^{2}}{8p}\Longrightarrow\nu_{i}\geq 1-\frac{1}{2\chi}(1-\lambda_{n})>\frac{(1-p)^{2}}{(1+p)^{2}}.$
As a result, when $\chi\geq\frac{1}{p}$, we have $\nu_{i}\in(c_{i},1)$, i.e., $\Delta_{i}(\nu_{i},p)<0$. It implies that
@@
$Q^{\mathrm{s}}_{2,i}=\left[\begin{array}[]{cc}\frac{1}{2}(p-1)\sqrt{1-\nu_{i}}+\frac{1}{2}j\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}&\frac{1}{2}(p-1)\sqrt{1-\nu_{i}}-\frac{1}{2}j\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}\\ p\sqrt{\nicefrac{{1}}{{2\chi}}}&p\sqrt{\nicefrac{{1}}{{2\chi}}}\end{array}\right]$
$(Q^{\mathrm{s}}_{2,i})^{-1}=\frac{\sqrt{2\chi}}{p\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}}\left[\begin{array}[]{cc}-jp\sqrt{\nicefrac{{1}}{{2\chi}}}&\frac{1}{2}\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}+\frac{1}{2}j(p-1)\sqrt{1-\nu_{i}}\\ jp\sqrt{\nicefrac{{1}}{{2\chi}}}&\frac{1}{2}\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}-\frac{1}{2}j(p-1)\sqrt{1-\nu_{i}}\end{array}\right]$
Since the spectral radius of matrix is upper bounded by any of its norm, $10<p_{0}\leq p<$ 1, and $10<\nu_{i}<$ 1, it holds that

### v4

@@ XV Proof of Theorem 5
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+p(1-p))}{\chi^{2}},$
@@
$\alpha\leq\min\left\{\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L},\sqrt{\frac{p(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{p^{3}(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},\ \chi\geq\frac{288(1-p)}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}\leq\frac{1+\gamma}{2}<1.$
Since $\alpha\leq\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L}$ and $\frac{\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}\leq\frac{8\chi^{2}}{p^{2}(1-\lambda_{2})^{2}}$, we have $\frac{\alpha^{2}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}\leq\frac{1}{384L^{2}}$. Thus, it holds that
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{n\alpha^{2}L^{2}}{96}\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{n\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2p(1-p)))}{192\chi^{2}}.$
@@
$\left[\begin{array}[]{c}\mathbb{E}\!\left[\big\|\bar{{\bf{x}}}^{t+1}-{\bf{x}}^{\star}\big\|^{2}\right]\\ \frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\right]\\ \end{array}\right]\leq\underbrace{\left[\begin{array}[]{cc}1-\mu\alpha&6\alpha L\\ \frac{\alpha^{2}L^{2}}{96}&\frac{1+\gamma}{2}\\ \end{array}\right]}_{:=A}\left[\begin{array}[]{c}\mathbb{E}\!\left[\big\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big\|^{2}\right]\\ \frac{1}{n}\mathbb{E}\!\left[\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}\right]\\ \end{array}\right]+\underbrace{\left[\begin{array}[]{c}\frac{\alpha^{2}\sigma^{2}}{n}\\ \frac{\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2p(1-p)))}{192\chi^{2}}\\ \end{array}\right]}_{:=b}.$ (78)
@@
$=\frac{1}{\mu\alpha(1-\gamma)(\frac{1}{2}-\frac{\alpha^{3}L^{3}}{16\mu(1-\gamma)})}\left[\begin{array}[]{cc}\frac{1-\gamma}{2}&6\alpha L\\ \frac{\alpha^{2}L^{2}}{96}&\mu\alpha\\ \end{array}\right]\left[\begin{array}[]{c}\frac{\alpha^{2}\sigma^{2}}{n}\\ \frac{\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2p(1-p)))}{192\chi^{2}}\\ \end{array}\right]$
$\leq\frac{4}{\alpha\mu(1-\gamma)}\left[\begin{array}[]{c}\frac{(1-\gamma)\alpha^{2}\sigma^{2}}{2n}+\frac{6L\alpha^{3}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2p(1-p)))}{192\chi^{2}}\\ \frac{\alpha^{4}L^{2}\sigma^{2}}{96n}+\frac{\mu\alpha^{3}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2p(1-p)))}{192\chi^{2}}\\ \end{array}\right],$
@@
$\|(I-A)^{-1}b\|_{1}\leq\frac{2\alpha\sigma^{2}}{\mu n}+\frac{(6L\alpha^{2}\sigma^{2}+\mu\alpha^{2}\sigma^{2})(192\chi^{2}+(4\chi^{2}+2p(1-p)))}{48\mu(1-\gamma)\chi^{2}}.$
@@
$\mathbb{E}\!\left[\big\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\big\|^{2}\right]\leq\Big(1-\frac{\alpha\mu}{4}\Big)^{t}a_{0}+\frac{2\alpha\sigma^{2}}{\mu n}+\frac{7L\alpha^{2}\sigma^{2}(192\chi^{2}+(4\chi^{2}+2p(1-p)))}{48\mu(1-\gamma)\chi^{2}}.$
Since ${\bf{X}}^{0}=[{\bf{x}}^{0},\cdots,{\bf{x}}^{0}]^{\sf T}$, by [42, (75)], we have $\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq 2\alpha^{2}\|({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-1}\|\|\nabla F({\bf{X}}^{0})-{\bf{1}}_{n}\otimes(\nabla f({\bf{x}}^{0}))^{\sf T}\|^{2}=\frac{2n\alpha^{2}\varsigma_{0}^{2}}{1-\gamma}$. Note that $\frac{1}{1-\gamma}\leq\frac{4\chi}{p(1-\lambda_{2})}$. It holds that
$a_{0}=\big\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big\|^{2}+\frac{1}{n}\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq\big\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big\|^{2}+\frac{8\chi\alpha^{2}\varsigma_{0}^{2}}{p(1-\lambda_{2})}$
@@ XVI Proof of Corollary 5
$c_{0}=\big\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big\|^{2},\ b_{0}=\frac{8\chi\varsigma_{0}^{2}}{p(1-\lambda_{2})},\ c_{1}=\frac{2\sigma^{2}}{\mu n},\ c_{2}=\frac{7L\sigma^{2}(192\chi^{2}+(4\chi^{2}+2p(1-p)))}{12\mu p(1-\lambda_{2})\chi}.$
@@
$\alpha\leq\frac{1}{\underline{\alpha}}\triangleq\min\left\{\frac{1}{2L},\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L},\sqrt{\frac{p(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\frac{72\mu}{L^{2}},\frac{1-\gamma}{12L+\nicefrac{{\mu}}{{2}}},\sqrt[3]{4\mu(1-\gamma)}\frac{1}{L}\right\}=\mathcal{O}\left(\frac{\mu p(1-\lambda_{2})}{\chi L^{2}}\right).$
@@
Collecting these cases together into (XVI), we have
@@ XVII Proof of Lemma 5
$\widetilde{{\bf{X}}}^{t+1}$ $=\widehat{{\bf{W}}}_{a}\widetilde{{\bf{Z}}}^{t}-{\bf{W}}_{b}{\bf{E}}^{t},$ (81b)
$\widetilde{{\bf{U}}}^{t+1}$ $=\widetilde{{\bf{U}}}^{t}+\frac{p^{2}}{2\chi}{\bf{W}}_{b}\widetilde{{\bf{Z}}}^{t}+p{\bf{E}}^{t}.$ (81c)
@@
$\left[\begin{array}[]{c}\widetilde{{\bf{X}}}^{t+1}\\ \widetilde{{\bf{U}}}^{t+1}\\ \end{array}\right]=\left[\begin{array}[]{cc}\widehat{{\bf{W}}}_{a}&-\widehat{{\bf{W}}}_{a}{\bf{W}}_{b}\\ \frac{p^{2}}{2\chi}{\bf{W}}_{b}&{\bf{I}}-\frac{p^{2}}{2\chi}{\bf{W}}_{b}^{2}\\ \end{array}\right]\left[\begin{array}[]{c}\widetilde{{\bf{X}}}^{t}\\ \widetilde{{\bf{U}}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\widehat{{\bf{W}}}_{a}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p^{2}}{2\chi}{\bf{W}}_{b}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]+\left[\begin{array}[]{c}-{\bf{W}}_{b}{\bf{E}}^{t}\\ p{\bf{E}}^{t}\\ \end{array}\right].$
@@
$\left[\begin{array}[]{c}{\bf{P}}^{-1}\widetilde{{\bf{X}}}^{t+1}\\ {\bf{P}}^{-1}\widetilde{{\bf{U}}}^{t+1}\\ \end{array}\right]=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{\Lambda}}}_{b}\\ \frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}&{\bf{I}}-\frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]\left[\begin{array}[]{c}{\bf{P}}^{-1}\widetilde{{\bf{X}}}^{t}\\ {\bf{P}}^{-1}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}{\bf{P}}^{-1}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}{\bf{P}}^{-1}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]+\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}{\bf{P}}^{-1}{\bf{E}}^{t}\\ p{\bf{P}}^{-1}{\bf{E}}^{t}\\ \end{array}\right].$
@@
$\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t+1}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t+1}\\ \end{array}\right]$ $=\left[\begin{array}[]{cc}\hat{{\bf{\Lambda}}}_{a}&-\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{\Lambda}}}_{b}\\ \frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}&{\bf{I}}-\frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]\left[\begin{array}[]{c}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{X}}}^{t}\\ \hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{U}}}^{t}\\ \end{array}\right]-\alpha\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]+\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ p\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right].$
@@
[…]
@@
$H^{\mathrm{s}}_{i}=\left[\begin{array}[]{cc}1-\frac{p}{2\chi}(1-\lambda_{i})&-(1-\frac{p}{2\chi}(1-\lambda_{i}))\sqrt{1-\lambda_{i}}\\ \frac{p^{2}}{2\chi}\sqrt{1-\lambda_{i}}&1-\frac{p^{2}}{2\chi}(1-\lambda_{i})\\ \end{array}\right].$
Setting $\nu_{i}=1-\frac{p}{2\chi}(1-\lambda_{i})$, we have $\nu_{i}\in(0,1)$ and $H_{i}$ can be rewritten as
$H^{\mathrm{s}}_{i}=\left[\begin{array}[]{cc}\nu_{i}&-\nu_{i}\sqrt{\frac{2\chi}{p}(1-\nu_{i})}\\ \frac{p^{2}}{2\chi}\sqrt{\frac{2\chi}{p}(1-\nu_{i})}&1-p(1-\nu_{i})\\ \end{array}\right].$
@@
Since $\nu_{i}=1-\frac{p}{2\chi}(1-\lambda_{i})\geq 1-\frac{p}{2\chi}(1-\lambda_{n}),i=2,\ldots,n$, $\chi\geq 1$ and $\lambda_{n}\in(-1,1)$, it holds that
$\nu_{i}\geq 1-\frac{p}{2\chi}(1-\lambda_{n})>\frac{(1-p)^{2}}{(1+p)^{2}},\ \forall i\in[n].$
As a result, we have $\nu_{i}\in(c_{i},1)$, i.e., $\Delta_{i}(\nu_{i},p)<0$. It implies that
@@
$Q^{\mathrm{s}}_{2,i}=\left[\begin{array}[]{cc}\frac{1}{2}(p-1)\sqrt{1-\nu_{i}}+\frac{1}{2}j\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}&\frac{1}{2}(p-1)\sqrt{1-\nu_{i}}-\frac{1}{2}j\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}\\ \sqrt{\nicefrac{{1}}{{2\chi}}}&\sqrt{\nicefrac{{1}}{{2\chi}}}\end{array}\right]$
$(Q^{\mathrm{s}}_{2,i})^{-1}=\frac{\sqrt{2\chi}}{\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}}\left[\begin{array}[]{cc}-j\sqrt{\nicefrac{{1}}{{2\chi}}}&\frac{1}{2}\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}+\frac{1}{2}j(p-1)\sqrt{1-\nu_{i}}\\ j\sqrt{\nicefrac{{1}}{{2\chi}}}&\frac{1}{2}\sqrt{(1+p)^{2}(\nu_{i}-c_{i})}-\frac{1}{2}j(p-1)\sqrt{1-\nu_{i}}\end{array}\right]$
Since the spectral radius of matrix is upper bounded by any of its norm, $0<p<1$, and $0<\nu_{i}<1$, it holds that